# Hypergraph clustering using Ricci curvature: an edge transport perspective

### v1

@@ 3 Two complementary notions of Ricci curvature on hypergraphs
  • Computing transport distances between measures defined on the nodes of a hypergraph typically relies on the clique expansion (see Section 2.1 and Hayashi et al. (2020)). As a consequence, this approach introduces a loss of structural information, as clique expansions fail to capture certain higher-order interactions.
The first challenge is straightforward to treat, for instance by taking the average or the maximum of all pairwise curvatures of nodes in a hyperedge (as in Coupette et al. (2022)). We will discuss different possibilities in Section 3.1. The second challenge constitutes the main contribution of this paper and will further be developed in Section 3.2 where we explore a method to transport hyperedges.
Note that we restrict our analysis to the Ollivier-Ricci curvature. Another notion, the Forman-Ricci curvature (Forman, 2003), has been extended to the case of hypergraphs in Leal et al. (2021). We refer to Samal et al. (2018) for a detailed comparison of these two curvature types in the graph case.
@@ 3.1 Nodes transport on the clique graph
Let $H$ be a hypergraph with nodes $V=(v_{1},\ldots,v_{n})$ and edges $E=(e_{1},\ldots,e_{m})$. To compute the Ricci curvature between two nodes $x$ and $y$, we can still define measures $\nu_{x}$ and $\nu_{y}$ on $x$ and $y$ and their neighbors following Equation 1, as the definition of a neighbor is unchanged. The first solution is simply to replace the hypergraph with its clique graph. Three possibilities are considered to obtain a clique graph from a hypergraph. For the sake of clarity, we assume the hypergraph to be unweighted. Let $A_{C}=(a_{ij})_{1\leq i,j\leq n}$ be the adjacency matrix of the clique expansion of $H$.
@@
    Another option is to sum the contributions of each hyperedge containing two adjacent points. We first define a similarity adjacency matrix $\tilde{A}_{i,j}\in\mathbb{R}^{n\times n}$ such that
    $\tilde{a}_{ij}=\sum_{e\in E}\omega(e)\mathds{1}_{(v_{i},v_{j})\in e},$
    where $\omega$ is a weight function. A usual choice is to take $\omega(e)=1/|e|$ to avoid overemphasizing large hyperedges. The final adjacency matrix $A$ is obtained by inverting the non-zero entries of $\tilde{A}$.
  3. (iii)
    A final option is proposed by Zhou et al. (2022) and relies on a Jaccard index weighting. Recall that for two finite sets $U$ and $V$, the Jaccard index $J(U,V)=\frac{|U\cap V|}{|U\cup V|}$ is a measure of similarity between sets $U$ and $V$. Let $H^{\star}=(V^{\star},E^{\star})$ be the dual hypergraph as defined in 2.1. We remark that for two nodes $v_{i}$ and $v_{j}$, $a_{ij}=0$ if and only if $|v^{\star}_{i}\cap v^{\star}_{j}|=0$. The weight of two adjacent nodes in the clique graph is given by $1/J(v^{\star}_{i},v^{\star}_{j})$. The Jaccard index is inverted to provide a dissimilarity weighting.
@@
The clustering pipeline given by Algorithm 1 then operates similarly by trimming all hyperedges with a Ricci flow larger than some threshold. The optimal threshold can be found by either maximizing the modularity of the clique graph, or a corresponding notion of hypergraph modularity, see Kamiński et al. (2024).
@@ 3.2 Edges transport on the line graph
As discussed in 2.1, replacing a hypergraph by its clique graph induces a substantial loss of information. To address this, we propose a novel notion of Ricci curvature specific to hypergraphs. Instead of transporting measures defined on the nodes of a hypergraph, we transport measures defined on its edges.
@@
The proposed method also computes pairwise curvature and then extends it to hyperedges using an aggregation step similar to the one previously defined. We first describe the Ricci curvature computation for a hypergraph $H=(V,E)$ with dissimilarity weights $d:E\to\mathbb{R}_{>0}$. Let $x$ and $y$ be two adjacent nodes in the hypergraph. We define probability measures $\mu_{x}$ and $\mu_{y}$ on their stars $St(x)$ and $St(y)$, see Figure 3.
@@
The most common choice is to consider $\mu_{x}$ and $\mu_{y}$ to be uniform probability measures. Similarly to what is commonly done for standard Ricci curvature on graphs, another option is given a parameter $\alpha\in[0,1]$, to define the following measure:
$\mu_{x}(e)=\begin{cases}\frac{\alpha d(e)}{C_{1}}&\text{if }x,y\in e\\ \frac{(1-\alpha)d(e)}{C_{2}}&\text{if }e\in St(x)\text{ and }e\notin St(y)\\ 0&\text{ otherwise, }\end{cases}$ (3)
where $C_{1}=\sum_{e\in St(x)\cap St(y)}d(e)$ and $C_{2}=\sum_{e\in St(x)\setminus St(y)}d(e)$ are normalization constants to ensure $\mu_{x}$ is a probability measure. The definition of $\mu_{y}$ is analogous. The distance $d_{H}$ between two edges $e$ and $f$ is defined using the line graph: $d_{H}(e,f)=d_{LG}(e^{\star},f^{\star})$ where $d_{LG}$ is the standard shortest path distance for graphs. Note that similarly to the clique graph, there are several ways the line graph can be weighted (for instance using intersection or Jaccard weights between edges). This in turn ensures that similarly to the graph case, we can define an optimal transport plan and a corresponding Wasserstein distance $W$ between the two measures $\mu_{x}$ and $\mu_{y}$. The edge-Ricci curvature $\kappa_{\mathcal{E}}$ between two nodes $x$ and $y$ is therefore defined as
@@
Pairs of nodes in different communities tend to have very different stars, which implies in turn a low curvature. Conversely, if the two nodes $x$ and $y$ belong to the same community, the measures $\mu_{x}$ and $\mu_{y}$ can be transported onto each other with a low cost, resulting in a curvature close to 1. We remark that even in the case of a 2-uniform hypergraph, this definition is not equivalent to the usual curvature on graphs from Section 2.2. In addition, we lose the interpretability observed in Figure 2 concerning the sign of the curvature.
@@
Similarly to the graph case, the curvature is then turned into a flow. We start with an initial edge weighting $w_{\mathcal{E}}^{(0)}=d$. At the $l$-th iteration, the weights are updated according to the following three steps:
@@
Similarly to the graph case this procedure is iterated $N$ times, and hyperedges with a weight larger than a given threshold are trimmed. The resulting connected components form the assigned node communities.
In traditional Ricci flow, negatively curved edges gradually gain weight, increasing the cost of moving probability measures through them. Here, we take a dual perspective: edges with low curvature are assigned increasing mass, making them progressively more expensive to transport. The line graph, however, remains unchanged, such that distances between edges are unaffected by the flow process.
@@ 3.3 A toy clustering example
We start by comparing the node-Ricci flow from Section 3.1 with the edge-Ricci flow from Section 3.2 in a synthetic example where transport maps and curvatures can be computed explicitly. We adopt a very similar setting to the one of Ni et al. (2019) where we consider a family of hypergraphs $H(a,b)$ where $a,b\geq 3$ are integers. Consider $b$ complete binary graphs $C_{1},\ldots,C_{b}$ on $a$ distinct vertices. For each complete graph $C_{i}$, take a particular vertex $v_{i}$ (called gateway vertex) and consider the hyperedge of size $b$, $(v_{1},\ldots v_{b})$ connecting all gateway vertices across the communities. This hypergraph exhibits a clear community structure where each $C_{i}$ forms a distinct community, and there is a single hyperedge connecting all the communities together. Figure 4 provides a visual representation. The case of the node-Ricci flow for an unweighted clique expansion has been treated explicitly in Ni et al. (2019), for measures defined by Equation 1 with $\alpha=0$. They observe that by symmetry, in the clique expansion, the Ricci flow at the $l-$th iteration can only take three possible values: $w_{\mathcal{N},1}^{(l)}$ for edges between two gateway nodes, $w_{\mathcal{N},2}^{(l)}$ for edges between a gateway node and another node from its community, and $w_{\mathcal{N},3}^{(l)}$ for edges between two non-gateway nodes from the same community. They further demonstrate that if $a>b$, $w_{\mathcal{N},3}^{(l)}=\left(\frac{1}{a}\right)^{l}\underset{l\to\infty}{\to}0$ and that there exists $\lambda>\frac{1}{a}$ and constants $c_{1}>c_{2}$ such that $w_{\mathcal{N},1}^{(l)}=c_{1}\lambda^{l}+o(\lambda^{l})$ and $w_{\mathcal{N},2}^{(l)}=c_{2}\lambda^{l}+o(\lambda^{l})$ as $l\to\infty$. This implies that the Ricci flow-based trimming algorithm properly identifies the communities $C_{1},\ldots,C_{b}$. In the case of the edge-Ricci flow, we have the following result:
@@
Keeping the same notation as above and considering the edge-Ricci flow with parameter $\alpha=0$ and aggregation $Agg$ with the maximum function. At the $l$-th iteration, the Ricci flow can only take three possible values: $w_{\mathcal{N},1}^{(l)}$ for the edge of size $b$ connecting all gateway nodes, $w_{\mathcal{N},2}^{(l)}$ for binary edges between a gateway node and another node from its community, and $w_{\mathcal{N},3}^{(l)}$ for binary edges between two non-gateway nodes from the same community. Under these circumstances, we have that for all $l>0$:
@@
The proof is deferred to Appendix A. Proposition 1 implies that the edge-Ricci flow can also identify the community structure with an edge cut-off parameter between 1 and 2. We notice that in this case, the explicit expression of the Ricci flow, although much simpler than in the node-Ricci case, does not have an exponential decay to 0 for the intra-community edges. This implies that the cut-off parameter might be more delicate to tune in practice. However, note that Proposition 1 holds for any values of $a$ and $b$, contrary to the result of Ni et al. (2019) which only holds for $a>b$. This implies that in practice, edge-Ricci allows for better detection of many small communities.
@@ 4 Experiments
To evaluate the practical effectiveness of Ricci flow for hypergraph clustering, we conducted a range of experiments on synthetic and real data. We remarked that the choice of the weights in the clique graph and the line graph had little impact on the final accuracy, and opted for a Jaccard type weighting. Curvatures of hyperedges are computed by aggregating curvatures of pairs of nodes using the maximum function. For node-Ricci flow, we have used the python library GraphRicciCurvature with default parameters for the measures $\nu$ defined on the nodes. We did not observe a great sensibility to the $\alpha$ parameter for the measures $\mu$ defined on the edges. The clustering accuracy is measured using the Normalized Mutual Information score (NMI). The implementations of edge-Ricci and node-Ricci flows for hypergraphs are available at https://github.com/OlympioH/Ricci_curv_hyp and are directly adapted from the GraphRicciCurvature library.
@@ 4.1 Hypergraphs stochastic block models
We first evaluate our method on a synthetic set-up to highlight differences between edges-Ricci and nodes-Ricci flows. We consider a hypergraph stochastic block model inspired by Kim et al. (2018). We consider a hypergraph with $n$ nodes and $k$ communities. We give ourselves two size parameters $s_{in}$ and $s_{out}$ and generate $N_{in}$ hyperedges of size $s_{in}$ between nodes of the same community for each community and $N_{out}$ hyperedges of size $s_{out}$ containing at least one node of each community. In the graph case where $s_{in}=s_{out}=2$, this is a form of stochastic block model, see Abbe (2018). This is a simple model where intra-community edges can be interpreted as signal and inter-community edges as noise. In the hypergraph case, as soon as $s_{in}\geq 3$, inter-community edges are also regrouping nodes of the same community together and cannot simply be interpreted as pure noise.
We try to reconstruct the communities using edges-Ricci and nodes-Ricci flow clustering algorithms. We consider randomly generated hypergraphs of $n=100$ nodes, $k=2$ equal-sized communities and compare the performances of the two methods for various values of $s_{in}$ and $s_{out}$. We fix the value of $N_{in}$ and measure the evolution of the clustering accuracy for a growing number $N_{out}$ of ”noisy” hyperedges. We compute $N=20$ flow iterations for each method and take the thresholding parameter $\tau$ that maximizes the NMI. We average over 10 random hypergraph generations and report the corresponding NMI in Figure 5. We remark that nodes-flow seems to perform comparatively better than edges-flow when the size of intra-community edges $s_{in}$ is large and the size of inter-community edges $s_{out}$ is small. More precisely, based on the numerical results of Figure 5, we can make the heuristic observation that edges-Ricci yields a better clustering performance than nodes-Ricci whenever $s_{in}<s_{out}-1$ and that nodes-Ricci performs better whenever $s_{in}\geq s_{out}$. When $s_{out}=s_{in}+1$, the two methods have a comparative performance, depending on the level of noise. These differences can be simply understood heuristically: when contaminating a hypergraph by adding a hyperedge of size $s$, the clique graph will have up to $s(s-1)/2$ new edges while the line graph will only have a single new additional node. This accounts for an improved robustness of edges-flow to large hyperedges contamination, as it mostly relies on shortest-path computations on the line-graph. The performance of edges-flow seems to be widely independent of the actual values of $s_{in}$ and $s_{out}$ and solely depends on the number of edges $N_{in}$ and $N_{out}$.

### v2

@@ 3 Two complementary notions of Ricci curvature on hypergraphs
  • Computing transport distances between measures defined on the nodes of a hypergraph typically relies on the clique expansion (see Section 2.1 and Hayashi et al., (2020)). As a consequence, this approach introduces a loss of structural information, as clique expansions fail to capture certain higher-order interactions.
The first challenge is straightforward to treat, for instance by taking the average or the maximum of all pairwise curvatures of nodes in a hyperedge (as in Coupette et al., (2022)). We will discuss different possibilities in Section 3.1. The second challenge constitutes the main contribution of this paper and will further be developed in Section 3.2 where we explore a method to transport hyperedges.
Note that we restrict our analysis to the Ollivier-Ricci curvature. Another notion, the Forman-Ricci curvature (Forman,, 2003), has been extended to the case of hypergraphs in Leal et al., (2021). We refer to Samal et al., (2018) for a detailed comparison of these two curvature types in the graph case.
@@ 3.1 Nodes transport on the clique graph
Let $H$ be a hypergraph with nodes $V=(v_{1},\ldots,v_{n})$ and edges $E=(e_{1},\ldots,e_{m})$. To compute the Ricci curvature between two nodes $x$ and $y$, we can still define measures $\nu_{x}$ and $\nu_{y}$ on $x$ and $y$ and their neighbors following Equation 1, as the definition of a neighbor is unchanged. We first propose to replace the hypergraph with its clique expansion. Two possibilities are considered to obtain a clique graph from a hypergraph. For the sake of clarity, we assume the hypergraph to be unweighted. Let $A_{C}=(a_{ij})_{1\leq i,j\leq n}$ be the adjacency matrix of the clique expansion of $H$.
@@
    The second option is proposed by Zhou et al., (2022) and relies on a Jaccard index weighting. Recall that for two finite sets $U$ and $V$, the Jaccard index $J(U,V)=\frac{|U\cap V|}{|U\cup V|}$ is a measure of similarity between sets $U$ and $V$. Let $H^{\star}=(V^{\star},E^{\star})$ be the dual hypergraph as defined in 2.1. We remark that for two nodes $v_{i}$ and $v_{j}$, $a_{ij}=0$ if and only if $|v^{\star}_{i}\cap v^{\star}_{j}|=0$. The weight of two adjacent nodes in the clique graph is given by $1/J(v^{\star}_{i},v^{\star}_{j})$. The Jaccard index is inverted to provide a dissimilarity weighting.
    A discussion on the relative performance of these two weightings can be found in Section 4.3.
@@
The clustering pipeline given by Algorithm 1 then operates similarly by trimming all hyperedges with a Ricci flow larger than some threshold. The optimal threshold can be found by either maximizing the modularity of the clique graph, or a corresponding notion of hypergraph modularity, see Kamiński et al., (2024).
@@ 3.2 Edges transport on the line graph
As discussed in 2.1, replacing a hypergraph by its clique graph induces a substantial loss of information. To address this, we propose a novel notion of Ricci curvature specific to hypergraphs. Instead of transporting measures defined on the nodes of a hypergraph, we transport measures defined on its edges. The motivation is that two nodes from different communities should be contained in sets of hyperedges that typically do not intersect much.
@@
The proposed method also computes pairwise curvature and then extends it to hyperedges using an aggregation step similar to the one previously defined. We first describe the Ricci curvature computation for a hypergraph $H=(V,E)$ with initial dissimilarity weights $d:E\to\mathbb{R}_{>0}$. Let $x$ and $y$ be two adjacent nodes in the hypergraph. We define simple probability measures $\mu_{x}$ and $\mu_{y}$ on their stars $St(x)$ and $St(y)$:
@@
$\mu_{x}(e)=\begin{cases}\frac{d(e)}{C}&\text{if }e\in St(x)\\ 0&\text{ otherwise, }\end{cases}$ (3)
where $C=\sum_{e\in St(x)}d(e)$ is a normalization constant to ensure that $\mu_{x}$ is a probability measure. The definition of $\mu_{y}$ is analogous. Similarly to the measures $\nu_{x}^{\alpha,p}$ defined on the nodes, the measures $\mu_{x}$ could be adapted to include hyperparameters $\alpha$ and $p$. However, preliminary experiments on a wide variety of settings suggested that it had close to no effect on the resulting curvature.
The distance $d_{H}$ between two edges $e$ and $f$ is defined using the line graph: $d_{H}(e,f)=d_{\mathbf{L}(G)}(e^{\star},f^{\star})$ where $\mathbf{L}(G)$ is the line graph of $H$, see Figure 3. Note that similarly to the clique graph, there are several ways the line graph can be weighted (for instance using Jaccard weights between hyperedges). This aspect will be discussed in more details in Section 4.3. Similarly to the graph case, we can then define an optimal transport plan and a corresponding 1-Wasserstein distance $W$ using the distance $d_{H}$ between the two measures $\mu_{x}$ and $\mu_{y}$. The edge-Ricci curvature $\kappa_{\mathcal{E}}$ between two adjacent nodes $x$ and $y$ is therefore defined as
@@
Pairs of nodes in different communities tend to have very different stars, which implies in turn a low curvature. Conversely, if the two nodes $x$ and $y$ belong to the same community, the measures $\mu_{x}$ and $\mu_{y}$ can be transported onto each other with a low cost, resulting in a curvature close to 1. We remark that even in the case of a 2-uniform hypergraph, this definition is not equivalent to the usual curvature on graphs from Section 2.2. In addition, we lose the interpretability observed in Figure 2 concerning the sign of the curvature. Note that although it is inspired by the notion of curvature developed by Ollivier, (2007), edge-Ricci curvature cannot be directly connected to Ollivier’s original definition of curvature in metric spaces. As such, deriving theoretical results for this curvature seems currently out of reach.
As edge-Ricci curvature uses both the clique and the line expansions, it allows to be more sensitive to the hypergraph information than node-Ricci curvature which erases information by considering the clique. For example, we can observe that pairs of nodes in Figure 1 mastrenghtsy have different edge-Ricci curvature while every pair has the same node-Ricci curvature. However, a general theoretical analysis of the differences between node-Ricci and edge-Ricci curvatures remains out of reach. In Section 3.3, we investigate these differences on a toy example and illustrate the strengths of each notion.
@@
Similarly to the graph case, the curvature is then turned into a discrete flow. We start with an initial edge weighting $w_{\mathcal{E}}^{(0)}=d$. At the $l$-th iteration, the weights are updated according to the following three steps:
@@
Similarly to the graph case this procedure is iterated $N$ times, and hyperedges with a weight larger than a given threshold are trimmed. The resulting connected components form the assigned node communities. In traditional Ricci flow with $p=0$, negatively curved edges gradually gain weight, increasing the cost of moving probability measures through them. Here, we take a dual perspective: edges with low curvature are assigned increasing mass, making them progressively more expensive to transport. The line graph, however, remains unchanged, such that distances between edges are unaffected by the flow process. We refer to Tian et al., (2025) for another example of method using the line graph for community detection.
@@ 3.3 A toy clustering example
We start by comparing the node-Ricci flow from Section 3.1 with the edge-Ricci flow from Section 3.2 in a synthetic example where transport maps and curvatures can be computed explicitly. We adopt a very similar setting to the one of Ni et al., (2019) where we consider a family of hypergraphs $H(a,b)$ where $a,b\geq 3$ are integers. Consider $b$ complete graphs $C_{1},\ldots,C_{b}$ on $a$ distinct vertices. For each complete graph $C_{i}$, take a particular vertex $v_{i}$ (called gateway vertex) and consider the hyperedge of size $b$, $(v_{1},\ldots v_{b})$ connecting all gateway vertices across the communities. This hypergraph exhibits a clear community structure where each $C_{i}$ forms a distinct community, and there is a single hyperedge connecting all the communities together. Figure 4 provides a visual representation.
The case of the node-Ricci flow for an unweighted clique expansion has been treated explicitly in Ni et al., (2019), for measures defined by Equation 1 with $\alpha=0$ and $p=0$. They observe that by symmetry, in the clique expansion, the Ricci flow at the $l$-th iteration can only take three possible values: $w_{\mathcal{N},1}^{(l)}$ for edges between two gateway nodes, $w_{\mathcal{N},2}^{(l)}$ for edges between a gateway node and another node from its community, and $w_{\mathcal{N},3}^{(l)}$ for edges between two non-gateway nodes from the same community. They further demonstrate that if $a>b$, $w_{\mathcal{N},3}^{(l)}=\left(\frac{1}{a}\right)^{l}\underset{l\to\infty}{\to}0$ and that there exists $\lambda>\frac{1}{a}$ and constants $c_{1}>c_{2}$ such that $w_{\mathcal{N},1}^{(l)}=c_{1}\lambda^{l}+o(\lambda^{l})$ and $w_{\mathcal{N},2}^{(l)}=c_{2}\lambda^{l}+o(\lambda^{l})$ as $l\to\infty$. This implies that the Ricci flow-based trimming algorithm properly identifies the communities $C_{1},\ldots,C_{b}$.
We derive a similar result for specific measures $\mu$. Equation 3 is slightly modified such that the transported measures are now defined, for two nodes $x$ and $y$ as:
[…]
where $C$ is a normalization constant. In practice, this choice of measure yields results extremely similar to the ones defined by Equation 3. It simply discards edges shared between $x$ and $y$, which would typically not play a big role in deriving the transport plan. This choice simply ensures a much simpler expression for the Ricci flow to facilitate the delivery of the following proposition:
@@
We keep the same notation as above and consider the edge-Ricci flow with measures $\tilde{\mu}$ defined by Equation 4 and aggregation $Agg$ with the maximum function. At the $l$-th iteration, the Ricci flow can only take three possible values: $w_{\mathcal{N},1}^{(l)}$ for the edge of size $b$ connecting all gateway nodes, $w_{\mathcal{N},2}^{(l)}$ for binary edges between a gateway node and another node from its community, and $w_{\mathcal{N},3}^{(l)}$ for binary edges between two non-gateway nodes from the same community. Under these circumstances, we have that for all $l>0$:
@@
The proof is deferred to Appendix A. Proposition 1 implies that the edge-Ricci flow can also identify the community structure with an edge cut-off parameter between 1 and 2. We notice that in this case, the explicit expression of the Ricci flow, although much simpler than in the node-Ricci case, does not have an exponential decay to 0 for the intra-community edges. This implies that the cut-off parameter might be more delicate to tune in practice. However, note that Proposition 1 holds for any values of $a$ and $b$, contrary to the result of Ni et al., (2019) which only holds for $a>b$. This implies that in practice, edge-Ricci allows for better detection of many small communities.
@@ 4 Experiments
To evaluate the practical effectiveness of Ricci flow for hypergraph clustering, we conducted a range of experiments on synthetic and real data. We opted for a Jaccard type weighting for the clique graph and the line graph. Curvatures of hyperedges are computed by aggregating curvatures of pairs of nodes using the maximum function. For node-Ricci flow, we have used the python library GraphRicciCurvature with parameters $p=1$ and $\alpha=0.5$ for the measures $(\nu_{x}^{\alpha,p})_{x\in V}$ defined on the nodes. We refer to Section 4.3 for a discussion about the choice of default values for the hyperparameters. The clustering accuracy is measured using the Normalized Mutual Information score (NMI). The implementations of edge-Ricci and node-Ricci flows for hypergraphs are available at https://github.com/AnonRicciHG/Ricci_curv_HG and are directly adapted from the GraphRicciCurvature library.
@@ 4.1 Hypergraphs stochastic block models
We first evaluate our method on a synthetic set-up to highlight differences between edge-Ricci and node-Ricci flows. We consider a hypergraph stochastic block model inspired by Kim et al., (2018). We consider a hypergraph with $n$ nodes and $k$ communities. We give ourselves two size parameters $s_{in}$ and $s_{out}$. We generate $N_{in}$ hyperedges of size $s_{in}$ by uniformly sampling $s_{in}$ nodes of the same community for each community. We generate $N_{out}$ hyperedges of size $s_{out}$ by first sampling one node from each community at random, and drawing the $s_{out}-2$ remaining nodes uniformly from $N$. In the graph case where $s_{in}=s_{out}=2$, this is a form of stochastic block model, see Abbe, (2018). This is a simple model where intra-community edges can be interpreted as signal and inter-community edges as noise. In the hypergraph case, as soon as $s_{in}\geq 3$, inter-community edges are also regrouping nodes of the same community together and cannot simply be interpreted as pure noise.
We try to reconstruct the communities using edge-Ricci and node-Ricci flows clustering algorithms. We consider randomly generated hypergraphs of $n=100$ nodes, $k=2$ equal-sized communities and compare the performances of the two methods for various values of $s_{in}$ and $s_{out}$. We fix the value of $N_{in}$ and measure the evolution of the clustering accuracy for a growing number $N_{out}$ of "noisy" hyperedges. We compute $N=20$ flow iterations for each method and take the thresholding parameter $\tau^{\star}$ that maximizes the NMI. We average over 5 random hypergraph generations and report the corresponding NMI in Figure 5. We remark that node-Ricci flow seems to perform comparatively better than edges-flow when the size of intra-community edges $s_{in}$ is large and the size of inter-community edges $s_{out}$ is small and conversely. These differences can be simply understood heuristically: when contaminating a hypergraph by adding a hyperedge of size $s$, the clique graph will have up to s(s1)/2s(s-1)/2 new edges while the line graph will only have a single new additional node. This accounts for an improved robustness of edges-flow to large hyperedges contamination, as it mostly relies on shortest-path computations on the line-graph.